OpenStudy (anonymous):

At the start of each week, May put $400 into her cheque account. During the week she would then withdraw $300. Assuuming that there are no bank or government charges, in which week would she have $1000 iin the bank for the first time?

OpenStudy (king):

10th week

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

(400 - 300 ) w = $1,000 100 w = $1,000 ->w = 1,000/ 100 = 10 wks ( week 10th)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Do you have any question?

OpenStudy (anonymous):

i don't think so..